OpenStudy (anonymous):

1. If I know the lengths of two sides of a right triangle, how do I find the third? 2. Could I find the two missing side lengths of a right triangle if I only know one side length and one angle measure (other than the 90 degree angle)? 3. Could I find the two missing angle measures if I know some of the side lengths of a right triangle? 4. What makes a triangle a “special” right triangle? How can special right triangles help me find side lengths? number the questions you are answering please

OpenStudy (anonymous):

A^2+B^2=C^2

OpenStudy (anonymous):

1. Use the pythagorean theorem. a squared plus b squared equals c squared. The hypotenuse needs to be c. 2. Yes, use sin, cos, or tan. 3. Yes, same as #2. 4. Special right triangles are 30, 60, 90 degree triangles or 45, 45, 90 degree triangles. There are special ratios between the side lengths of these two types which you can use to solve for the missing side length.

OpenStudy (anonymous):

Can you give some examples? and also elaborate with 2 and 3

OpenStudy (anonymous):

Are you working with numbers or just theory?

OpenStudy (anonymous):

define theory in that context?

OpenStudy (anonymous):

Are you solving problems (calculating answers) or are you just describing in words the answers for any numbers you might be given?

OpenStudy (anonymous):

i have to do this : Create a unique teaching example for each question. Work through each example step by step, and use complete sentences to help your friend with the questions.

OpenStudy (anonymous):

i just dont know how to explain that with sin cos and tan

OpenStudy (anonymous):

1. Suppose you have two sides lengths of a right triangle where one leg is 3 and the hypotenuse is 5. Using the pythagorean theorem, plug 3 into A and 5 into C. Solve for leg b. 2. If you know one of the angle measures of a right triangle and one of the side lengths, you can set up a trignometrical equation to solve for the other side lengths. "Standing" at the known angle, figure out if you know the side across from you (the opposite side), the side next to you (the adjacent side) or the hypotenuse. Also figure out which one of these sides you want to find. Figuring out what two sides you are dealing with, you can figure out which trig function to use. Opposite side / hypotenuse is the SIN, Adjacent side/ hypotenuse is the COS, and opposite side / adjacent side is the TAN. Next, set up your equation. ex. Sin 32 degrees = 5/x (this means I have the opposite side and am trying to find the hypotenuse's length)

OpenStudy (anonymous):

Repeat again for the other side you want to find OR use the Pythagorean Theorem now that you know two sides :) 3. Use the same formula but this time you won't know the angle measure. You will know both of the side lengths for the right side of the equation. After dividing the right side, take the inverse of the trig function you are trying to find. EX. cos X = 1/2, cos X = 0.5, so the cos to the -1 of 0.5 equals the angle measure of X!

OpenStudy (anonymous):

I'm working on my graduate homework so read that and let me know if I can be of anymore help... and I'll check back in a minute...

OpenStudy (anonymous):

That helps so much,thanks so much<3

OpenStudy (anonymous):

no problem! I wish I could work on that stuff instead of matrix theory! Are you in school to be a teacher?

OpenStudy (anonymous):

no, its just part of an assignment, i have no idea why and what class do you learn matrix theory in?

OpenStudy (anonymous):

the class is "Matrix Theory" haha it is the graduate class after Linear Algebra.

OpenStudy (anonymous):

you're in college?

OpenStudy (anonymous):

yeah, and im a high school math teacher... you?

OpenStudy (anonymous):

i'm a junior in highschool

OpenStudy (anonymous):

Okay, well I teach seniors mostly. Glad I could help!